Hi all, I was hoping to get help with the following problems. I need to find ∆y / ∆x, given the following information:
a) y = 1/x, x = 3, ∆x = 0.01;
b) y = sqrt(x), x = 9, ∆x = 0.9
Any help would be great,
Thanks, -nick.

Answer by Alan3354(69443)

I need to find ∆y / ∆x, given the following information:
a) y = 1/x, x = 3, ∆x = 0.01;
@ x = 2.99, y = 1/2.99, ∆y= 1/3 - 1/2.99 = (2.99 - 3)/8.97 =~ -0.0011148
@ x = 3, y= 1/3
@ x = 3.01, y = 1/3.01. Vy= 1/3.01 - 1/3 = (3-3.01)/9.03 =~ -0.0011074
------------
b) y = sqrt(x), x = 9, ∆x = 0.9
Excel is very useful for these.